# I Introduction

### v1

@@ III.1 Electron energy loss spectrum and secondary electron yield
Using the MC approach described above, we have calculated the reflection electron energy loss spectrum (REELS) [33] for different initial energies of the incident electrons (10, 20, 30, 50 keV), as shown in Fig. (5). Part of the electrons of the primary beam can be backscattered with the same energy as the incident energy and represents the so-called elastic (or lossless) peak, which typically also hides substructures due to electron-phonon scattering. To obtain a good signal-to-noise ratio, we performed all MC calculations with a total number of electrons equal to .
@@
The elastic peak is extremely bright in the figure and lies at the initial kinetic energy of the primary bream (black solid line for 10 keV kinetic energy, red line for 20 keV, green line for 30 keV and blue line for 50 keV). The energy loss spectra show several peaks reflecting the ELF lineshape. In Fig. (5), the inset image highlights the first 100 eV of energy loss for an initial impact energy of 20 keV. The shoulders in the 20 keV REELS at , and eV from the elastic peak are directly related to single-particle excitations (Br 4 states Pb 6 states and Br 4 states Cs 6 states), which also occur in the ELF. The highest peak is at eV away from the elastic peak, which is a mixed plasmon.
To investigate the effects of the film thickness on the reflected electrons, we also calculated the backscattering coefficient (i.e. the fraction of backscattered electrons to total electrons) as a function of incident electron energies of , , , and keV. In Figure (6) we show the results for three CsPbBr3 thin films of thickness (cyan coloured line with full circles), (black coloured line with full squares) and nm (magenta coloured line with full diamonds). The backscattering coefficients for the CsPbBr3 sample show a trend characterised by a relative minimum and a maximum that does not shift with increasing thickness. However, the number of reflected electrons increases with thickness and reaches a similar value for all thicknesses at an initial kinetic energy of the primary beam of 50 keV. The elastic peaks in Fig. (5) have an intensity trend similar to the coefficient. This highlights the decisive contribution of elastic collision in reflection. To simulate perovskite samples that may be charged, we performed the MC calculations for a 100 nm thick layer for three different WF: 4.1 eV (i.e. the experimental value of the WF for the pristine material [34, 35]), 3 eV and 5 eV. The MC calculations for different WFs on a 100 nm slab (black, dotted and dashed lines) overlap almost perfectly with the results for the 4.1 eV WF and confirm that its value does not significantly influence the backscattering of the primary electrons.
@@
Electron beams that hit a material can trigger the emission of secondary electrons by depositing their kinetic energy in the target. These electrons are emitted by the atoms of the solid as a result of inelastic interactions with the target, which may involve either the primary beam or other secondary electrons. The emission of secondary electrons is typically influenced by individual and collective charges, Auger decay, plasmons, phonons, trapping and charge density waves. Secondary electron emission is the primary contrast mechanism in an SEM and in SE imaging within the TEM. Secondary electron generation can also contribute to the charging of a sample which has implications on electron beam-induced damage mechanisms [4, 7, 5, 36]. In this context, the SE yield is defined as the ratio between the number of secondary electrons (reflected or transferred) and the total number of primary electrons. The secondary yield is thus defined per incident particle.
@@
In Fig. (7) we show the yield of reflected secondary electrons determined for different initial energies of the primary electron beam of keV for thicknesses of 50 (cyan-coloured line), 100 (black lines) and 200 nm (solid magenta-coloured line) and WF = 4.1 eV (pristine material). To investigate the dependence of the yield on the WF, calculations were also performed for WF = 3 eV (black dash-dotted line) and 5 eV (black dashed line) and 100 nm thickness.
@@
The reflection yield of secondary electrons decreases with increasing initial kinetic energy, from which we can conclude that the low-energy primary electrons have a greater chance of producing the escaping secondary electrons immediately below the target surface. The secondary production dependent on the low energy impinging electron is confirmed by the similar trend independent of the sample thickness. In contrast to the analysis of primary electrons, the WF plays a fundamental role in increasing the production of secondary electrons: At low kinetic energy of the primary beam, a WF = 3 eV leads to an increased production of secondary electrons, while a WF = 5 eV attenuates them. This is particularly evident at the lowest energies (e.g. 2 and 5 keV) and confirms that impinging electrons with lower kinetic energy have a higher probability of emitting secondary electrons. In contrast, the electrons penetrate deep into the sample at the high energy of the primary beam (50 keV) and the generated secondary electrons cannot escape at the vacuum-target interface due to the numerous inelastic scatterings they suffer on their way out of the solid; in this case, the different WFs have little influence on the secondary reflection yield.
@@
In Fig. (8) the secondary electron spectra are shown as a function of the initial kinetic energy of the primary beam equal to 1 (orange solid line), 2 (black solid line), 5 (red solid line), 10 (green solid line), 20 (blue solid line) and 50 keV (cyan solid line) for three different thicknesses of 50 (left), 100 (centre), 200 nm (right). The resulting secondary electrons are distributed over a range of 25 eV with a sustained peak of around 2.6 eV for all layer thicknesses and kinetic energies of the primary beam. The general shapes are consistent with the secondary kinetic energy spectra obtained by real-time TDDFT (RT-TDDFT) calculations for graphene, where a peak position at 3.5 eV [37] with a WF=4.56 eV was found, in semiconductor materials such as TiN, VN, GaAs, InAs, InSb, PbS by using Reverse MC approach [38], and with the analytical expression originally developed by Chung–Everhart [39] for metals, which relates the position of the secondary electron peak to the WF. The Chung–Everhart expression can be also used to estimate grossly the peak distributions for non-metals.
@@
Considering the CsPbBr3 WF = eV and using the Chung–Everhart formula, the peak position is eV ( 1.3 eV below our estimate). Nevertheless, in our simulations, we confirm a red (blue) shift of the peak position by 0.3 eV for a WF of 3.0 eV (5.0 eV). At a thickness of 100 nm (centre panel of Fig. 8), the positively charged sample (WF = 3 eV, dotted-dashed line) enhances the intensity of the spectra of the kinetic energy of the secondary electrons, while the WF=5 eV (corresponding to negatively charged samples, dashed line) attenuates it. The conclusions drawn for the reflected secondary electron yield are confirmed by these results: (i) The more intense secondary electron distributions are found for low kinetic energies of the primary electrons rather than for higher energies (e.g. 50 keV). (ii) The WFs influence the distributions of the secondary electron yield mainly at low kinetic energies of the primary electrons (with a maximum at keV). (iii) The distributions are not significantly influenced by the thickness of the layer.
III.2 Transmission of the primary and secondary electrons.
@@
The analysis carried out for the backscattered electrons is repeated for the primary and secondary electron transmission. In Fig. (9), the transmission coefficient () is plotted against the initial kinetic energies of the primary beam for three thicknesses. The number of transmitted primary electrons per incident particle is of course always less than 1 at low kinetic energy, as a large fraction of the incident electrons are reflected. Nevertheless, the primary transmission coefficient saturates at a higher kinetic energy (10 keV).
In the case of 50 nm (cyan solid line), the primary electrons are almost completely transmitted at an incident energy of 2 keV. After 10 keV, however, decreases slightly.
In the 100 nm thick layer (black solid line), the transmission coefficient of the primary electrons shows a behaviour between 50 (cyan solid line) and 200 nm (magenta solid line), where is reached at 10 keV kinetic energy. Like the backscattering coefficient of the primary beam, the WF also does not influence the electron transmission for primary electrons, as was demonstrated in particular for the 100 nm thick layer (small and large squares).
@@
The transmission coefficient of the secondary electrons as a function of the primary beam kinetic energy reported in Fig. (10) shows a maximum, whose position and intensity depend on the thickness: 2 keV for 50 nm (cyan solid line), 5 keV for 100 nm (black solid line) and 10 keV for 200 nm (magenta solid line).
@@
We also note that the intensity of the maximum of decreases with increasing thickness and the WF also significantly influences the intensity of the transmitted secondary electrons, the higher the WF, the smaller . We attribute this result to the fact that the only secondary electrons that can escape are those that originate near the vacuum interface (opposite the entrance surface). This behaviour can be interpreted in light of the Fig. (11), where we present the distribution of transmitted secondary electrons as a function of origin depth. Furthermore, Fig. (11) shows that the number of secondary electrons produced initially increases when the energy of the primary electrons increases. However, this trend is reversed as soon as the energy exceeds a certain threshold value, which is determined by the thickness of the film. This leads to a saturation point at which a considerable number of electrons from the primary beam are transmitted and can therefore no longer contribute to the production of secondary electrons. This is also reflected in the thickness-dependent plateau in Fig. (9) for the transmission of the primary electrons. Finally, we find that the maximum of secondary electron transmission occurs at a depth of about Å along the direction of incidence from the interface to the vacuum and the generation of secondary electrons is filtered out at low initial kinetic energy of the primary beam, e.g. the 2 keV distribution is most intense in a 50 nm thick layer, while it is attenuated in the 100 nm thick layer and suppressed in the 200 nm thick layer.
@@
Finally, Figure (12) shows the energy distributions of the transmitted secondary electrons for different layer thicknesses parameterised with respect to the initial kinetic energy of the primary beam. Similar to the behaviour of the reflected secondary electrons, the energy spectra of the transmitted secondary electrons show a persistent peak around 2.6 eV, independent of the thickness and kinetic energy of the primary beam, with a tail reaching up to 20 eV. WF of 3 (5) eV increases (attenuates) the intensity of the peaks and produces a small red (blue) shift of the peak position, as demonstrated in the centre. The WF impacts the transmission yield of secondary electrons, even at a kinetic energy of 50 keV.
@@ IV Discussion and conclusions
In this study, we theoretically investigated the scattering processes of high-energy primary and secondary electrons expected from Scanning Transmission Electron Microscopy (STEM) measurements of CsPbBr3. This was achieved through Monte Carlo (MC) simulations combined with first-principles LR-TDDFT calculations for Electron Loss Function (ELF).
For the first time, we calculated the LR-TDDFT ELF and the Inelastic Mean Free Path (IMFP) for bulk CsPbBr3 to support these simulations. We particularly focused on quantifying the reflection and transmission for backscattering and secondary electron yield (SEY) processes, as well as the energy distributions of the emitted secondary electrons.
Our findings showed that, for backscattering processes involving primary electrons, there is nearly no dependence on the WF, with a transmission coefficient approaching one at input energies of 10 keV. However, the situation changes significantly for the reflection and transmission of secondary electrons. The WF plays a critical role in the production of these outgoing electrons, although its influence diminishes at higher energies.
In terms of transmission, the observed peaks are dependent on both the thickness of the material and the initial kinetic energy of the primary electrons. Outgoing secondary electrons are most likely to be emitted when they are generated within the last 10 nm of the nanolayer. Therefore, the kinetic energy of the incoming electrons must be carefully balanced; it should be high enough to generate secondary electrons without releasing too much energy and low enough to prevent the generation of secondaries too deeply within the sample.
These results contribute to a better understanding of the transmission and reflection characteristics of primary and secondary electrons, ultimately aiding in the optimization of STEM experimental measurements.
@@ Acknowledgements.
This work was supported by the Ada Lovelace Centre.The authors acknowledge the use of the ARCHER2 (via the UK Car-Parrinello Consortium, EP/X035891/1) and STFC-SCARF HPC facilities.
The authors receiving funds from grant agreement No. 10104665 acknowledge the European Union.
P.E.T. thanks Dr J. Jackson for assistance with the Questaal calculations and Dr I. Scivetti for his help with Archer2.
Appendix A Drude-Lorentz method
The inelastic scattering cross-section is derived from the ELF in the entire spectrum of excitation energies () and momentum transfers . Performing ab initio calculations of the ELF over a wide energy range is often prohibitively expensive due to the complex computational requirements resulting from the inclusion of multiple electronic transitions to excited states. While first-principles methods can effectively calculate the ELF for energies below 100 eV, these calculations had to be extended to higher excitation energies by using experimental NIST X-ray atomic data (TDDFT-NIST) up to 400 keV [30].

### v2

@@ III.1 Electron energy loss spectrum
Part of the electrons of the primary beam can be backscattered with the same energy as the incident energy and represents the so-called elastic (or lossless) peak, which typically also hides substructures due to electron-phonon scattering. We performed all MC calculations with a total number of electrons equal to to obtain a good signal-to-noise ratio.
@@
The elastic peaks shown in Fig. 5 are extremely bright and lie at the initial kinetic energy of the primary beam. The energy loss spectra show several peaks reflecting the ELF lineshape. In Fig. 5, the inset image highlights the first 100 eV of energy loss for an initial impact energy of 20 keV. The shoulders in the 20 keV REELS at , and eV from the elastic peak are directly related to single-particle excitations (Br 4 states Pb 6 states and Br 4 states Cs 6 states), which also occur in the ELF. The highest peak is at eV away from the elastic peak, which is a mixed single-electron/plasmon transition.
III.2 Backscattered electrons
To investigate the effects of the film thickness on the reflected electrons, we also calculated the backscattering coefficient (i.e. the fraction of backscattered electrons to total electrons) as a function of incident electron energies of , , 15, , , 40 and keV.
In Fig. 6 we show the results for three CsPbBr3 thin films of thickness (cyan coloured line with full circles), (black coloured line with full squares) and nm (magenta coloured line with full diamonds). The backscattering coefficients for the CsPbBr3 sample exhibit a monotonically decreasing trend with increasing kinetic energy, which depends on the sample thickness; in particular, for a given initial kinetic energy of the primary electrons, the reflection is higher for thicker samples and is almost zero at 50 keV, i.e. all electrons are transmitted. At 5 keV, the reflection rate for the 200 nm sample is more than twice as high as for the 50 nm sample ( = 0.52 compared to = 0.22, respectively). The intensities of the elastic peaks in Fig. 5 show the decisive contribution of the elastic collisions to the backscattering coefficient.
III.3 Secondary electrons
@@
Electron beams that hit a material can also trigger the emission of secondary electrons by depositing their kinetic energy in the target. These electrons are emitted by the atoms of the solid as a result of inelastic interactions with the target, which may involve either the primary beam or other secondary electrons. The yield of transmitted or reflected secondary electrons recorded by the detector is influenced by all kinds of inelastic interactions with individual and collective charges, such as Auger decay, phonons, plasmons, trapping and charge density waves (CDW). Secondary electron emission is the primary contrast mechanism in an SEM and in SE imaging within the TEM. Secondary electron generation can also contribute to the charging of a sample, which has implications on electron beam-induced damage mechanisms [10, 13, 11, 49].
In this context, the SE yield is defined as the ratio between the number of secondary electrons (reflected or transmitted with a kinetic energy below 50 eV, although secondary electrons above this nominal limit can also be found) and the total number of primary electrons. The secondary yield is thus defined per incident particle.
@@
In Fig. 7 we show the yield of reflected secondary electrons determined for different initial kinetic energies of the primary electron beam of keV for thicknesses of 50 (cyan line), 100 (black lines) and 200 nm (magenta line) for EA = 4.1 eV (pristine material). To investigate the dependence of the yield on the EA, which may depend on surface termination and material history, calculations were also performed for EA = 3 eV (black dash-dotted line) and 5 eV (black dashed line) and 100 nm thickness.
@@
The reflection yield of secondary electrons decreases with increasing initial kinetic energy, from which we can conclude that low-energy primary electrons have a greater chance of producing the escaping secondary electrons immediately below the target surface. The secondary reflection at low kinetic energy of the incident electrons shows a behaviour that is almost independent of the sample thickness, in contrast to the backscattering coefficient of the reflected primary electrons. In addition, the EA plays a fundamental role in increasing the production of secondary electrons. At low kinetic energy of the primary beam, an EA = 3 eV leads to increased production of secondary electrons, while an EA = 5 eV attenuates them. This is particularly evident at the lowest energies (e.g. 2 and 5 keV) and confirms that impinging electrons with lower kinetic energy have a higher probability of emitting secondary electrons. At odds, the electrons penetrate deep into the sample at the high energy of the primary beam (50 keV) and the generated secondary electrons cannot escape at the vacuum-target interface due to the numerous inelastic scatterings they suffer on their way out of the solid; in this case, the different EAs have little influence on the secondary reflection yield.
@@
In Fig. 8 the secondary electron energy (, with the near-surface vacuum energy) spectra are shown as a function of the initial kinetic energy of the primary beam equal to 1 (orange solid line), 2 (black solid line), 5 (red solid line), 10 (green solid line), 20 (blue solid line) and 50 keV (cyan solid line) for three different thicknesses of 50 (left), 100 (centre), 200 nm (right). The resulting secondary electrons are distributed over a range of 25 eV with a sustained peak of around 2.6 eV for all layer thicknesses and kinetic energies of the primary beam. These lineshapes are consistent with the secondary kinetic energy spectra obtained by real-time TDDFT (RT-TDDFT) calculations for graphene, where a peak position at 3.5 eV [50] with an EA=4.56 eV was found, for semiconductor materials such as TiN, VN, GaAs, InAs, InSb, PbS using the reverse MC approach [51], and with the analytical expression for metals originally developed by Chung–Everhart [52], which relates the position of the secondary electron peak to the EA. The Chung–Everhart expression can also be used to estimate grossly the peak distributions for non-metals.
@@
Considering the CsPbBr3 EA = eV, the Chung–Everhart formula provides a peak position measured from the vacuum level at 1/3 of the EA, i.e. eV ( 1.3 eV below our estimate). Nevertheless, in our simulations, we confirm a red (blue) shift of the peak position by 0.3 eV for an EA of 3.0 eV (5.0 eV). At a thickness of 100 nm (centre panel of Fig. 8), the positively charged sample (EA = 3 eV, dotted-dashed line) enhances the intensity of the spectra of the kinetic energy of the secondary electrons, while the EA=5 eV (corresponding to negatively charged samples, dashed line) attenuates it.
The conclusions that can be drawn from these results for the reflected secondary electron yield are as follows: (i) The more intense secondary electron distributions are found for low kinetic energies (5-10 keV) of the primary electrons rather than for higher energies (e.g. 50 keV). (ii) The EAs influence the distributions of the secondary electron yield mainly at low kinetic energies of the primary electrons (with a maximum at keV). (iii) The distributions are not significantly influenced by the thickness of the layer.
III.4 Transmission of the primary and secondary electrons.
@@
The analysis carried out for the backscattered electrons is repeated for the primary and secondary electron transmission. In Fig. 9, the transmission coefficient () is plotted against the initial kinetic energies of the primary beam for three thicknesses. The number of transmitted primary electrons per incident particle is always less than 1 at low kinetic energy, as a large fraction of the incident electrons are reflected. However, the primary transmission coefficient saturates at a higher kinetic energy ( keV) for all investigated samples. We also note that the transmission coefficient decreases with increasing thickness at all energies. In particular, for the 100 nm thick layer (black solid line), the transmission coefficient of the primary electrons shows a behaviour between 50 (cyan solid line) and 200 nm (magenta solid line). In the case of 50 nm (cyan solid line), the primary electrons are almost completely transmitted at an incident energy of 5 keV. We emphasise that the spectra and yield of the secondary electrons, which are typically orders of magnitude greater in number than the primary electrons, have a better signal-to-noise ratio than the corresponding magnitudes of the primary electrons.
@@
The transmission coefficient of the secondary electrons as a function of the primary beam kinetic energy reported in Fig. 10 shows a maximum, whose position and intensity depend on the thickness: 2 keV for 50 nm (cyan solid line), 5 keV for 100 nm (black solid line) and 10 keV for 200 nm (magenta solid line).
@@
We also find that the intensity of the maximum of decreases with increasing thickness and the EA also significantly affects the intensity of the transmitted secondary electrons; in particular the higher the EA, the smaller . We attribute this result to the fact that the only secondary electrons that can escape are those that originate near the vacuum interface (opposite the entrance surface). This behaviour can be interpreted in the light of Fig. 11, where we plot the distribution of transmitted secondary electrons as a function of origin depth. Fig. 11 shows that the number of generated secondary electrons initially increases as the energy of the primary electrons increases. However, this trend reverses as soon as the energy exceeds a certain threshold value, which is determined by the thickness of the film. This leads to a saturation point at which a considerable number of electrons from the primary beam pass through the film without much interaction and no longer contribute to the production of secondary electrons. This is also reflected in the thickness-dependent plateau in Fig. 9 for the transmission of primary electrons ( 30 keV). Moreover, we find that the maximum secondary electron transmission occurs at a depth of about nm along the direction of incidence from the interface to the vacuum. Finally, we also note that the generation of secondary electrons is filtered out at a low initial kinetic energy of the primary beam, e.g. the 2 keV distribution is most intense in a 50 nm thick layer, while it is attenuated in the 100 nm thick layer and suppressed in the 200 nm thick layer.
@@
Finally, Fig. 12 shows the energy distributions of the transmitted secondary electrons for different parametrised layer thicknesses. Similar to the behaviour of the reflected secondary electrons, the energy spectra of the transmitted secondary electrons also exhibit a persistent peak around 2.6 eV, independent of the thickness and kinetic energy of the primary beam, with a tail reaching up to 20 eV. EA of 3 (5) eV increases (attenuates) the intensity of the peaks and produces a small red (blue) shift in the peak position, as shown in the centre panel of Fig. 12. The EA affects the transmission yield of secondary electrons, even at a kinetic energy of 50 keV.
@@ IV Discussion and conclusions
In this study, we have theoretically investigated the scattering processes of high-energy primary and secondary electrons expected from scanning transmission electron microscopy measurements of the perovskite CsPbBr3. To this end, we used MC simulations in combination with first-principles LR-TDDFT calculations of the electron loss function to determine the inelastic mean free path of bulk CsPbBr3.
We have focused in particular on analysing the reflection and transmission properties of backscattered and secondary electrons by calculating the spectra and yields of the emitted secondary electrons in thin films of CsPbBr3.
Our results show that the backscattering coefficient of the primary electrons decreases monotonically with kinetic energy and increases with film thickness, while the transmission coefficient is larger for thinner films and approaches unity for initial kinetic energies of the primary beam above 30 keV for all thicknesses investigated.
In the reflection and transmission of secondary electrons, the EA plays a critical role in producing these outgoing electrons, although its influence diminishes at higher energies.
In terms of transmission, the observed peaks depend on both the thickness of the material and the initial kinetic energy of the primary electrons. Outgoing secondary electrons are most likely to be emitted when they are generated within the last 10 nm of the nanolayer. Therefore, the kinetic energy of the incoming electrons must be carefully balanced; it should be high enough to generate secondary electrons without releasing too much energy and low enough to prevent the generation of secondary electrons too deeply within the sample.
These results contribute to a better understanding of the transmission and reflection characteristics of primary and secondary electrons, ultimately aiding in the quantitative interpretation of intensity in STEM experimental measurements.
@@ Acknowledgements.
This work was supported by the Ada Lovelace Centre. The authors gratefully acknowledge the use of ARCHER2 (via the UK Car-Parrinello Consortium, EP/X035891/1) and the STFC-SCARF HPC facilities. S.T. acknowledges the European Union for funding under grant agreement n° 10104665. P.E.T. thanks Dr J. Jackson for assistance with the Questaal calculations and Dr I. Scivetti for his help with Archer2.Appendix A Drude-Lorentz method
The inelastic scattering cross section is derived from the ELF in the entire spectrum of excitation energies () and momentum transfers . While first-principles methods can effectively determine the ELF for energies below 100 eV, performing ab-initio calculations of the ELF over such a broad energy range (up to 400 keV) is often prohibitively expensive due to the enormous number of excited states that must be included in the ELF calculation. To obtain the IMFP in the investigated energy range, we need to rely on experimental NIST X-ray atomic data (TDDFT-NIST) [43].